OpenStudy (anonymous):

I am learning quadratic formulas...I need help solving this equation...x(x-3)=x-9

OpenStudy (saifoo.khan):

x(x-3)=x-9 x^2 - 3x = x - 9 x^2 - 3x + x + 9 = 0 x ^2 - 2x + 9 = 0 for quadratic equation your equation must be equal to, ax^2 + bx + c. so here, a = 1 b = -2 c = 9. insert these values in the quadratic equation and solve. :)

OpenStudy (anonymous):

do i enter the a b and c in again after i have got the x^2-2x+9=0 or is that just the end of it